i need to know how to work out f(k)=7(3-k)+3k-4
And i need to find the solutions.
F ( k)= - 19
F ( k) = 37
F (k) = - 43

Answer by stanbon(75887)

i need to know how to work out
f(k)=7(3-k)+3k-4
And i need to find the solutions.
F ( k)= - 19
Solve
7(3-k)+3k-4 = -19
21-7k+3k-4 = -19
-4k +17 = -19
-4k = -36
k = 9
-------------------------------
F ( k) = 37
7(3-k)+3k-4 = 37
-4k+17 = 37
-4k = 20
k = -5
-------------------------------
F (k) = - 43
7(3-k)+3k-4 = 43
-4k+17 = 43
-4k = 26
k = -6 1/2
==================
Cheers,
Stan H.
===========